Determine the mean and standard deviation of the variable X in each of the following binomial distributions. a. n = 4 and a =Assume a Poisson distribution. a. If a = 2.5, find P(X = 1). c. If a = 0.5, find P(X = 4). b. If 1 = 8.0, find P(X = 5). d. IFrom a shipment of 25 cars being sent to local automobile dealers, 3 are SUVs. What is the probability that if 3 cars arriveThe dean of a business school wishes to form an executive committee of 3 from among the 50 tenured faculty members at the schAssume a Poisson distribution. Find the following probabilities. a. Let 2 = 6.0, find P(X23). b. Let 2 = 0.4, find P(XS1). c.
Answer #1

Question - Auswa Answer @ Mean = (4)(0-10) = 0.40 Standard deviatou = 744)(0-10)(0.90) - 10:36 = 0.60 - Answer D Mean = (5)(0